# Zero-shot OCR Accuracy of Low-Resourced Languages: A Comparative Analysis on Sinhala and Tamil

### v1

@@ 1 Introduction
While OCR systems have advanced significantly, particularly for High-Resourced Languages (HRL) such as English and French Nazeem et al. (2024), recognising text from complex or low-quality images, historical documents, and Low-Resource Languages (LRL) still presents challenges Agarwal and Anastasopoulos (2024). In this study, we conduct a thorough examination of various multilingual and monolingual OCR systems, assessing their capabilities for two selected low-resource languages in South Asia: Sinhala and Tamil.
@@
Sinhala is an Indo-European language spoken as L1 by just 16 million people mostly located in the island of Sri Lanka de Silva (2025). Sinhala has a script that is unique to it which descends from the Indian Brahmi Script Fernando (1949). Tamil is a Dravidian language spoken as L1 by around 79 million people located primarily in India, Sri Lanka, and Singapore Wijeratne et al. (2019). It also has a unique script that is also a descendant of the Indian Brahmi Script Paneerselvam (1972). Both Sinhala and Tamil are considered LRLs by the criteria proposed by Ranathunga and de Silva (2022), where Sinhala is deemed to be lower resourced (Category 02) as opposed to Tamil (Category 03).
@@ 2 Existing Works
Despite extensive research over the past several decades, the challenge of accurately recognising Sinhala characters in optical character recognition (OCR) systems remains a formidable obstacle Anuradha et al. (2020). Indic languages, such as Tamil, present a myriad of complexities and character variations, significantly complicating the development of effective OCR solutions. In clear terms, the accuracy of South Asian rounded scripts significantly trails behind that of Latin-based scripts, highlighting a crucial area for improvement and development Anuradha et al. (2021).
@@ 2.1 Sinhala OCR Systems
Several research studies have been conducted on developing Sinhala OCR systems. A system utilising the Tesseract 4.0 OCR engine with a graphical user interface has been proposed for the Sinhala language by Anuradha et al. (2020). The system comprises five main components: the user, an API, the Tesseract engine Smith (2007), a post-processor, and a data store. The Tesseract engine processes images and recognises text using LSTM-based deep learning techniques. However, some characters are not recognised by the Tesseract engine. To address this issue, the post-processor identifies those unrecognised characters and applies linguistic rules to ensure accurate output. For this study, commercially used font types were employed, varying in size, which resulted in an average accuracy of 94%.
@@
A study on multi-style printed Sinhala character recognition Maduranga and Jayalal (2022) utilised a hybrid Artificial Neural Network (ANN) model, incorporating concepts from previous research. The process includes four steps: Data Preprocessing for image enhancement and noise removal; Feature Extraction, dividing 50x50px character images into 9 zones with 12 pieces each to create a feature vector of 108 signals for the ANN; Development and Training, using line features from an 850-character database (primarily Iskoola Pota font [^2] https://learn.microsoft.com/en-us/typography/font-list/iskoola-pota) to train a back-propagation network in MATLAB with 108 input nodes, 78 hidden nodes, and 34 output nodes over 138 epochs, achieving around 75% training accuracy; and Testing, where a separate dataset of 1253 characters is used to evaluate performance.
@@
Velayuthan and Ambegoda (2025) conducted a comparative analysis addressing the challenge of digitising documents in low-resource languages by benchmarking several prominent OCR models: Surya [^3] https://github.com/VikParuchuri/surya, TR-OCR [^4] https://huggingface.co/Ransaka/TrOCR-Sinhala, EasyOCR [^5] https://github.com/JaidedAI/EasyOCR, and Tesseract OCR Smith (2007). While it initially states a focus on Sinhala and Tamil, the research ultimately conducted experiments and reported results only for Sinhala using two synthetic Sinhala datasets, as well as for English using the FUNSD dataset Jaume et al. (2019). Employing metrics such as CER and WER, the evaluation process involved data filtration and post-processing techniques. The findings revealed that Surya significantly outperformed the other models on the Sinhala datasets. Although all models demonstrated high error rates on the English dataset, Surya was identified as the most balanced option for Sinhala, exhibiting relatively high accuracy while maintaining moderate computational demands and superior power efficiency compared to TR-OCR. The discrepancy in accuracy between Sinhala and English can largely be attributed to the nature of the datasets, as the Sinhala datasets were synthetically generated, whereas the English dataset includes noisy images of form documents.
@@ 2.2 Tamil OCR Systems
Various research initiatives have been undertaken to create OCR systems for the Tamil language. Unlike Sinhala, which is primarily spoken in Sri Lanka, the Tamil language has a much wider geographical spread across the South Asian region. Despite their different linguistic roots, Tamil also utilises a rounded script similar to Sinhala due to their writing systems being related as depicted in Figure 1, which leads to similar challenges in the development of OCR technology.
Liyanage et al. (2015) developed a Tamil OCR system using the open-source Tesseract OCR engine, inspired by its applications with scripts such as Sinhala and Bangla. The methodology involved creating a 169-character OCR alphabet and preparing training data from selected words in various Unicode fonts. Researchers tested different training combinations and found that a model using data from three fonts at three sizes achieved the best results. The system was evaluated on 20 scanned images from ancient Tamil books, achieving an accuracy of 81%, which is a 12.5% improvement over Tesseract’s existing Tamil module.
Recent research introduced the Nayana framework Kolavi et al. (2025), which enhances Vision-Language Models (VLMs) such as GOT OCR Wei et al. (2024) for low-resource languages, including Tamil. It tackles data scarcity through a layout-aware synthetic data generation pipeline and Low-Rank Adaptation (LoRA). This system translates English documents into Tamil while retaining their layout, followed by a two-phase Cross-Modal Alignment training with LoRA. Nayana-OCR achieved notable improvements in performance, with a Word Error Rate (WER) of 0.551 and a METEOR score of 0.592, significantly outperforming the base GOT OCR model (WER 1.020, METEOR 0.051) and other traditional OCR systems, including Tesseract Smith (2007) and PaddleOCR [^6] https://github.com/PaddlePaddle/PaddleOCR, on the Tamil test set.
@@ 3 Methodology
As discussed earlier, researchers are utilising open-source tools to effectively fine-tune models for new languages and improve existing capabilities. Many of these tools offer multilingual support. Additionally, several organisations have developed commercial engines that excel in OCR. In this study, we thoroughly evaluate the capabilities of selected OCR engines for Sinhala and Tamil languages in a zero-shot setting.
@@ 3.1 Overview of Selected OCR Technologies
Cloud Vision API [^7] https://cloud.google.com/vision/docs/ocr: The Cloud Vision API enables developers to seamlessly incorporate vision detection features into their applications. This includes functionalities such as image labelling, face and landmark detection, OCR, and the tagging of explicit content. The first version of the API launched for general availability in May 2017. The API is designed to perform OCR on files like PDFs and TIFFs, as well as on images with dense text. It’s particularly optimised for image documents containing large amounts of text and those that feature handwriting, allowing for accurate recognition and conversion to machine-readable text.
@@
Tesseract: One of the most well-known and frequently used OCR engines is Tesseract OCR. This open-source project, originally developed by HP and now funded by Google, offers excellent text recognition capabilities. Tesseract combines Hidden Markov Models (HMMs) and various machine learning algorithms with traditional computer vision techniques to recognise text. The Tesseract 4.0 incorporates deep learning methods that significantly improve performance compared to earlier versions that primarily relied on conventional approaches. The deep learning models utilised by Tesseract are based on Long Short-Term Memory (LSTM) networks Smith (2007); Nazeem et al. (2024). In this study, we utilise Tesseract 5.5.0, an improved version of Tesseract 4.0, which incorporates the existing LSTM engine, along with several performance improvements.
@@
Subasa OCR [^9] https://ocr.subasa.lk/: The study by Anuradha et al. (2020) was extended, which involved a comprehensive series of deep learning (LSTM) based Tesseract 4.0 experiments to estimate the complexity of Sinhala OCR by investigating the effects of text genre, image resolution, and algorithmic complexity. Training data was prepared primarily from the UCSC 10M Sinhala corpus [^10] https://ltrl.ucsc.lk/tools-and-resourses/, using various Sinhala fonts and image qualities, involving character segmentation and iterative training processes. Evaluation was conducted on 30 non-identical test images categorised as old newspapers (200 DPI), old books (72 DPI), and contemporary books (300 DPI), with an additional test on low-DPI (96px) contemporary images. Accuracy was the primary metric, calculated by comparing character counts between the original text and the OCR output. Tesseract 4.0 achieved up to 67.02% character accuracy on old newspapers, up to 87.53% on old books, and up to 87.63% on contemporary books, even maintaining high accuracy up to 87.88% on low-DPI contemporary images Anuradha et al. (2021).
@@
Surya3: This is an OCR toolkit that supports over 90 languages and benchmarks favorably against cloud services. It features line-level text detection in any language, layout analysis (including detection of tables, images, headers, etc.), reading order detection, and table recognition (detecting rows and columns), as well as LaTeX OCR capabilities. The text detection model was trained on four A6000 GPUs for three days using a diverse set of images. It was built from the ground up using a modified EfficientViT architecture Liu et al. (2023) for semantic segmentation. Meanwhile, the text recognition model was trained on the same hardware for two weeks, utilising a modified Donut model Kim et al. (2022) that incorporates Grouped Query Attention (GQA) Ainslie et al. (2023), a Mixture of Experts (MoE) layer Shazeer et al. , UTF-16 decoding, and changes to the layer configuration. It is important to note that this system is designed for printed text and not for handwriting.
@@
EasyOCR5: This is an OCR technique that supports over 80 languages. EasyOCR utilises ResNet He et al. (2016), LSTM, and CTC (Connectionist Temporal Classification) Graves et al. (2006) models for character recognition. The detection component of EasyOCR employs the CRAFT Algorithm Baek et al. (2019). EasyOCR consists of three key elements. The first is feature extraction, which is executed by the ResNet model. The second element is sequence labelling, for which the LSTM algorithm is utilised, and the last component is decoding. Decoding relies on CTC. The Readtext function of EasyOCR is used during the recognition process. A notable feature of EasyOCR is its ability to read letters and numbers from images and provide the coordinates of their location Awalgaonkar et al. (2021).
@@
The Google Cloud Vision API and Document AI are commercial engines, whereas Google Tesseract, Surya, and EasyOCR are open-source systems that can be fine-tuned. Additionally, we selected Subasa OCR, which is a fine-tuned version of the Tesseract model available through a web application9, though the source code and model are not directly accessible.
@@ 3.2 Dataset Selection and Assembly
To achieve optimal results, we employed distinct datasets for each of the two selected languages, ensuring a tailored approach that enhances the effectiveness of our analysis. For the Sinhala language, we chose a dataset published in Hugging Face, sinhala_synthetic_ocr_large [^11] https://huggingface.co/datasets/Ransaka/sinhala_synthetic_ocr-large by Ravihara (2024), consisting of 6,969 pairs of images and reference texts.
@@
The Tamil text was obtained from OPUS [^12] https://opus.nlpl.eu/, specifically selecting OpenSubtitles v2024131313https://opus.nlpl.eu/OpenSubtitles/ta&en/v2024/OpenSubtitles. The content was then divided line by line, focusing solely on Tamil characters since the primary goal was language evaluation, resulting in 2,437,960 records. Subsequently, this set was filtered to retain only texts longer than 40 characters, yielding 222,658 records. This filtration step ensured that word-level evaluation was accurate. However, to ensure a fair comparison with the Sinhala dataset, we decided to randomly select 7,000 samples from the remaining texts to equalise the sample sizes. Regardless, conducting OCR on additional samples is challenging due to the resource consumption involved.
We carefully selected six unique fonts to diversify the creation of images from the text records, ensuring that each visual representation is impactful for the analysis. The selected fonts are as follows:
A function was then developed utilising the capabilities of the Pillow library to systematically convert textual data into image files. Its fundamental purpose is to ensure a proportional distribution of input text records across the defined set of font files, promoting an equitable use of fonts in the generated image dataset. For each text entry, the function dynamically calculates optimal image dimensions based on the measurements of the text bounding box, rendering the text in black on a white background. A significant feature of this implementation is the precise centring of the text within the generated image, achieved through calculated positioning in relation to the image’s height and width, thereby enhancing visual consistency and quality.
The postprocessing phase for the Tamil dataset involved the exclusion of records that contained no string values for both reference and generated features. In contrast, the postprocessing for the Sinhala dataset incorporated an additional step aimed at eliminating all non-Sinhala characters from both the reference and generated texts. This step was implemented to ensure that the evaluation concentrated exclusively on the linguistic capabilities of the system. Since we created the Tamil dataset, the postprocessing step focused on the language characters was addressed during preprocessing. This novel synthetic Tamil OCR public dataset is one of our contributions in this study1. Some examples of Tamil sentences from our synthetic dataset is shown in Figure 3.
Figure 3: Three examples of Tamil sentences from our data set in the fronts Hind Madurai, Anek Tamil, and Kavinar, respectively.

### v2

@@ 1 Introduction
While OCR systems have advanced significantly, particularly for High-Resourced Languages (HRL) such as English and French Nazeem et al. (2024), recognising text from complex or low-quality images, historical documents, and Low-Resource Languages (LRL) still presents challenges Agarwal and Anastasopoulos (2024). In this benchmarking study, we conduct a thorough examination of various multilingual and monolingual OCR systems, assessing their capabilities for two selected low-resource languages in South Asia: Sinhala and Tamil.
@@
Sinhala is an Indo-European language spoken as L1 by just 16 million people, mostly located in the island of Sri Lanka de Silva (2025). Sinhala has a script that is unique to it which descends from the Indian Brahmi Script Fernando (1949). Tamil is a Dravidian language spoken as L1 by around 79 million people located primarily in India, Sri Lanka, and Singapore Wijeratne et al. (2019). It also has a unique script that is also a descendant of the Indian Brahmi Script Paneerselvam (1972). Both Sinhala and Tamil are considered LRLs by the criteria proposed by Ranathunga and de Silva (2022), where Sinhala is deemed to be lower resourced (Category 02) as opposed to Tamil (Category 03).
@@ 2 Existing Works
Despite extensive research over the past several decades, the challenge of accurately recognising Sinhala characters in OCR systems remains a formidable obstacle Anuradha et al. (2020). Indic languages, such as Tamil, present a myriad of complexities and character variations, significantly complicating the development of effective OCR solutions. In clear terms, the accuracy of South Asian rounded scripts is significantly behind that of Latin-based scripts, highlighting a crucial area for improvement and development Anuradha et al. (2021).
@@ 2.1 Sinhala OCR Systems
Several research studies have been conducted on developing Sinhala OCR systems. A system utilising the Tesseract 4.0 OCR engine with a graphical user interface has been proposed for the Sinhala language by Anuradha et al. (2020). The system comprises five main components: the user, an Application Programming Interface (API), the Tesseract engine Smith (2007), a post-processor, and a data store. The Tesseract engine processes images and recognises text using Long Short Term Memory (LSTM) based deep learning techniques. However, some characters are not recognised by the Tesseract engine. To address this issue, the post-processor identifies those unrecognised characters and applies linguistic rules to ensure accurate output. For this study, commercially used font types were employed, varying in size, which resulted in an average accuracy of 94%.
@@
A study on multi-style printed Sinhala character recognition Maduranga and Jayalal (2022) utilised a hybrid Artificial Neural Network (ANN) model, incorporating concepts from previous research. The process includes four steps: Data Preprocessing for image enhancement and noise removal; Feature Extraction, dividing 50x50px character images into 9 zones with 12 pieces each to create a feature vector of 108 signals for the ANN; Development and Training, using line features from an 850-character database (primarily Iskoola Pota font [^2] https://learn.microsoft.com/en-us/typography/font-list/iskoola-pota) to train a backpropagation network in MATLAB with 108 input nodes, 78 hidden nodes, and 34 output nodes over 138 epochs, achieving around 75% training accuracy; and Testing, where a separate dataset of 1253 characters is used to evaluate performance.
@@
Velayuthan and Ambegoda (2025) conducted a comparative analysis addressing the challenge of digitising documents in low-resource languages by benchmarking several prominent OCR models: Surya [^3] https://github.com/VikParuchuri/surya, TR-OCR [^4] https://huggingface.co/Ransaka/TrOCR-Sinhala, EasyOCR [^5] https://github.com/JaidedAI/EasyOCR, and Tesseract OCR Smith (2007). While it initially states a focus on Sinhala and Tamil, the research ultimately conducted experiments and reported results only for Sinhala using two synthetic Sinhala datasets, as well as for English using the FUNSD dataset Jaume et al. (2019). The evaluation process, employing metrics such as CER (Character Error Rate) and WER (Word Error Rate), involved data filtration and post-processing techniques. The findings revealed that Surya significantly outperformed the other models on the Sinhala datasets. Although all models demonstrated high error rates on the English dataset, Surya was identified as the most balanced option for Sinhala, achieving relatively high accuracy while keeping moderate computational demands. Additionally, Surya demonstrated superior power efficiency, using an average of 0.69 kWh less across the two Sinhala datasets compared to TR-OCR. The accuracy difference between Sinhala and English is largely due to the datasets used; Sinhala datasets are synthetically generated, while the English dataset consists of noisy form document images.
@@ 2.2 Tamil OCR Systems
Various research initiatives have been undertaken to create OCR systems for the Tamil language. Unlike Sinhala, which is primarily spoken in Sri Lanka, the Tamil language has a much wider geographical spread across the South Asian region. Notwithstanding their different linguistic roots, Tamil also utilises a rounded script similar to Sinhala, as depicted in Figure 1, due to their writing systems being related. This leads to similar challenges in the development of OCR technology.
Liyanage et al. (2015) developed a Tamil OCR system using the open-source Tesseract OCR engine, inspired by its applications with scripts such as Sinhala and Bangla. The methodology involved creating a 169-character OCR alphabet and preparing training data from selected words in various Unicode fonts. They tested different training combinations and found that a model using data from three fonts at three sizes achieved the best results. The system was evaluated using 20 scanned images from ancient Tamil books, achieving an accuracy of 81%. This was benchmarked only against Tesseract’s existing Tamil module, over which, a 12.5% improvement was shown.
Recent research introduced the Nayana framework Kolavi et al. (2025), which enhances Vision-Language Models (VLMs) such as GOT (General OCR Theory) Wei et al. (2024) for low-resource languages, including Tamil. It tackles data scarcity through a layout-aware synthetic data generation pipeline and Low-Rank Adaptation (LoRA) Hu et al. (2022). This system translates English documents into Tamil while retaining their layout, followed by a two-phase Cross-Modal Alignment training with LoRA. Nayana-OCR achieved notable improvements in performance, with a Word Error Rate (WER) of 0.551 and a Metric for Evaluation of Translation with Explicit ORdering (METEOR) score of 0.592, significantly outperforming the base GOT model (WER 1.020, METEOR 0.051) and other traditional OCR systems, including Tesseract Smith (2007) and PaddleOCR [^6] https://github.com/PaddlePaddle/PaddleOCR, on the Tamil test set.
@@ 3 Methodology
As discussed earlier, contemporary studies are utilising open-source tools to effectively fine-tune models for new languages and improve existing capabilities. Many of these tools offer multilingual support. Additionally, several organisations have developed commercial engines that excel in OCR. In this study, we thoroughly evaluate the capabilities of selected OCR engines for Sinhala and Tamil languages in a zero-shot setting.
@@ 3.1 Overview of Selected OCR Technologies
Cloud Vision API [^7] https://cloud.google.com/vision/docs/ocr: The Cloud Vision API enables developers to seamlessly incorporate vision detection features into their applications. This includes functionalities such as image labelling, face and landmark detection, OCR, and the tagging of explicit content. The first version of the API launched for general availability in May 2017. The API is designed to perform OCR on files such as Portable Document Formats (PDFs) and Tag Image File Formats (TIFFs), as well as on images with dense text. It is particularly optimised for image documents containing large amounts of text and those that feature handwriting, allowing for accurate recognition and conversion to machine-readable text.
@@
Tesseract: One of the most well-known and frequently used OCR engines is Tesseract OCR. This open-source project, originally developed by Hewlett-Packard (HP) and now funded by Google, offers excellent text recognition capabilities. Tesseract combines Hidden Markov Models (HMMs) and various machine learning algorithms with traditional computer vision techniques to recognise text. The Tesseract 4.0 incorporates deep learning methods that significantly improve performance compared to earlier versions that primarily relied on conventional approaches. The deep learning models utilised by Tesseract are based on LSTM networks Smith (2007); Nazeem et al. (2024). In this study, we utilise Tesseract 5.5.0, an improved version of Tesseract 4.0, which incorporates the existing LSTM engine, along with several performance improvements.
@@
Subasa OCR [^9] https://ocr.subasa.lk/: The study by Anuradha et al. (2020) was extended, which involved a comprehensive series of deep learning (LSTM) based Tesseract 4.0 experiments to estimate the complexity of Sinhala OCR by investigating the effects of text genre, image resolution, and algorithmic complexity. Training data was prepared primarily from the UCSC 10M Sinhala corpus [^10] https://ltrl.ucsc.lk/tools-and-resourses/, using various Sinhala fonts and image qualities, involving character segmentation and iterative training processes. Evaluation was conducted on 30 non-identical test images categorised as old newspapers (200 DPI), old books (72 DPI), and contemporary books (300 DPI), with an additional test on low-DPI (96px) contemporary images. Accuracy was the primary metric, calculated by comparing character counts between the original text and the OCR output. This OCR engine achieved up to 67.02% character accuracy on old newspapers, up to 87.53% on old books, and up to 87.63% on contemporary books, even maintaining high accuracy up to 87.88% on low-DPI contemporary images Anuradha et al. (2021).
@@
Surya3: This is an OCR toolkit that supports over 90 languages and benchmarks favourably against cloud services. It features line-level text detection in any language, layout analysis (including detection of tables, images, headers, etc.), reading order detection, and table recognition (detecting rows and columns), as well as LaTeX OCR capabilities. The text detection model was trained on four A6000 GPUs for three days using a diverse set of images. It was built from the ground up using a modified EfficientViT architecture Liu et al. (2023) for semantic segmentation. Meanwhile, the text recognition model was trained on the same hardware for two weeks, utilising a modified Donut model Kim et al. (2022) that incorporates Grouped Query Attention (GQA) Ainslie et al. (2023), a Mixture of Experts (MoE) layer Shazeer et al. (2017), Unicode Transformation Format-16 (UTF-16) decoding, and changes to the layer configuration. It is important to note that this system is designed for printed text and not for handwriting.
@@
EasyOCR5: This is an OCR technique that supports over 80 languages. EasyOCR utilises ResNet He et al. (2016), LSTM, and Connectionist Temporal Classification (CTC) Graves et al. (2006) models for character recognition. The detection component of EasyOCR employs the Character Region Awareness For Text detection (CRAFT) Algorithm Baek et al. (2019). EasyOCR consists of three key elements. The first is feature extraction, which is executed by the ResNet model. The second element is sequence labelling, for which the LSTM algorithm is utilised, and the last component is decoding which relies on CTC. The EasyOCR’s Readtext function is utilized during text recognition which can read letters and numbers from images while providing their location coordinates Awalgaonkar et al. (2021).
@@
The Google Cloud Vision API and Document AI are commercial engines, whereas Google Tesseract, Surya, and EasyOCR are open-source systems. Additionally, we selected Subasa OCR, which is a fine-tuned version of the Tesseract model available through a web application9, though the source code and model are not directly accessible.
@@ 3.2 Dataset Selection and Assembly
As noted by Ranathunga et al. (2024); de Silva (2025), free and open datasets for this task are scarce for low-resourced languages; even when some published research may exist. To achieve optimal results, we employed distinct datasets for each of the two selected languages, ensuring a tailored approach that enhances the effectiveness of our analysis. For the Sinhala language, we chose a dataset published in Hugging Face, sinhala_synthetic_ocr-large [^11] https://huggingface.co/datasets/Ransaka/sinhala_synthetic_ocr-large by Ravihara (2024), consisting of 6,969 pairs of images and reference texts created using five different font families;
@@
The Tamil text was obtained from OPUS [^17] https://opus.nlpl.eu/ Tiedemann (2012), specifically selecting OpenSubtitles Lison and Tiedemann (2016) v2024181818https://opus.nlpl.eu/OpenSubtitles/ta&en/v2024/OpenSubtitles. The content was then divided line by line, focusing solely on Tamil characters since the primary goal was language evaluation, resulting in 2,437,960 records. Subsequently, this set was filtered to retain only texts longer than 40 characters, yielding 222,658 records. This filtration step ensured that word-level evaluation was accurate. However, to ensure a fair comparison with the Sinhala dataset, we decided to randomly select 7,000 samples from the remaining texts to equalise the sample sizes. This step was necessary to prevent evaluation scores from being skewed by differences in dataset size. Further, conducting OCR on additional samples is challenging due to the resource consumption involved.
We carefully selected six unique font families out of a total of 17 from Google Fonts [^19] https://fonts.google.com/ to diversify the creation of images from the text records, ensuring that each visual representation is impactful for the analysis. We identified the unique fonts visually, as some fonts, such as Mukta Malar [^20] https://fonts.google.com/specimen/Mukta+Malar and Baloo Thambi 2212121https://fonts.google.com/specimen/Baloo+Thambi+2, have characters that can appear very similar to the naked eye. The selected fonts are as follows:
A function was then developed utilising the capabilities of the Pillow [^28] https://pillow.readthedocs.io/ library to systematically convert textual data into image files. Its fundamental purpose is to ensure a proportional distribution of input text records across the defined set of font files, promoting an equitable use of fonts in the generated image dataset. For each text entry, the function dynamically calculates optimal image dimensions based on the measurements of the text bounding box, rendering the text in black on a white background. A significant feature of this implementation is the precise centring of the text within the generated image, achieved through calculated positioning in relation to the height and width of the image, thereby enhancing visual consistency and quality.
The post-processing phase for the Tamil dataset involved the exclusion of records that contained no string values for both reference and generated features. In contrast, the post-processing for the Sinhala dataset incorporated an additional step aimed at eliminating all non-Sinhala characters from both the reference and generated texts. This step was implemented to ensure that the evaluation concentrated exclusively on the targeted linguistic capabilities of the system. Since we created the Tamil dataset, the postprocessing step focused on the language characters was addressed during preprocessing. This novel synthetic Tamil OCR public dataset is one of our contributions in this study1. Some examples of Tamil sentences from our synthetic dataset are shown in Figure 3.
Figure 3: Three examples of Tamil sentences from our dataset in the fonts Hind Madurai, Anek Tamil, and Kavinar, respectively.